Write the equation of the line that passes through (5,8), and (3,2)

thank you

Answer by Alan3354(69443)

Write the equation of the line that passes through (5,8), and (3,2)
----------
Method 1 - the long way
1st, find the slope, m
m = diffy/diffx
m = (8-2)/(5-3) = 6/2
m = 3
Then
y-y1 = m*(x-x1) where (x1,y1) is the point
y-2 = 3*(x-3)
y-2 = 3x-9
y = 3x-7 (slope-intercept form)
3x-y = 7 (standard form)
----------
Method 2
|x y 1|
|5 8 1| = 0
|3 2 1|
------
x*(8-2) - y*(5-3) + (10-24) = 0
6x - 2y = 14
3x - y = 7